OpenStudy (anonymous):

Solve |2y - 2| = y. {2/3, 2} {1/3, 2}

OpenStudy (campbell_st):

2y - 2 = y and 2y - 2 = -y y = 2 3y = 2 y = 2/3

OpenStudy (anonymous):

or 2, 5

OpenStudy (campbell_st):

lol.. you would have 10 - 2 = 5 so 5 doesn't work